• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name dhamilton

    Log entry time 08:10:37 on April11,2008

    Entry number 231350

    keyword=Owl Shift Summary

    Owl Shift Summary
    Pi0 production kinematics D

    D. Hamilton
    B. Hahn
    K. Adrashev

    00:00 - shift started with 3 microamps on target during run 2771.

    00:40 - ROC5 reboot.

    01:20 - rebooted iocha5 (BCM), altered coefficients.

    02:05 - lost beam due to a problem with an arc vacuum valve.

    06:30 - beam restored to hall, data-taking resumed following a BB HV trip.

    07:10 - beam taken away while MCC try to fix problems with 5-pass beam for other halls.

    07:25 - beam restored to hall.

    08:00 - shift ends.